# Existence, uniqueness, and regularity of

### v1

@@ 5.2 Proof of Theorem 1
By Theorem 1.2.2.2 in Grisvard, (2011), both $\mathcal{X}$ and $\mathcal{X}_{n}$, $n\in\mathbb{N}$, satisfy a uniform exterior cone condition of size $(L,\theta)$, for some $L,\theta>0$ independent of $n$. In particular $(L,\theta)$ depends only on $M$ and $diam(\mathcal{X})$.
@@
$\begin{cases}\max\left\{u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,g-u\right\}=0\text{ in }\mathcal{Y}_{n}\\ u=b^{n}\text{ on }\partial\mathcal{Y}_{n},\end{cases}$ (Eqn)
@@
By Lemma 3, for all $n\in\mathbb{N}$, there exists a unique $L^{p}$-solution $u^{n}\in W^{1,2,p}\left(\mathcal{Y}_{n}\right)\subset W^{1,2,p}_{loc}\left(\mathcal{Y}_{n}\right)\cap\mathcal{C}^{0}\left(\bar{\mathcal{Y}}_{n}\right)$ of (Eqn). Extending all the the $u^{n}$’s by $u^{n}(x,t)=b(x,t)$ on $\bar{\mathcal{Y}}\setminus\mathcal{Y}_{n}$, we obtain a sequence $\left(u^{n}\right)_{n\in\mathbb{N}}\subset W^{1,2,p}\left(\mathcal{Y}\right)\subset W^{1,2,p}_{loc}\left(\mathcal{Y}\right)\cap\mathcal{C}^{0}\left(\bar{\mathcal{Y}}\right)$.
@@
We start with (i). By Lemma 4 with $D>T\vee diam(\mathcal{X})$ (using Lemma 12.1.9 in Krylov, (2018) to uniformly control $\left\|u^{n}\right\|_{L^{\infty}\left(\mathcal{Y}\right)}$, $n\in\mathbb{N}$), the sequence $\left(u^{n}\right)_{n\in\mathbb{N}}$ is equicontinuous. As a result, it has a convergent subsequence in $\mathcal{C}^{0}\left(\bar{\mathcal{Y}}\right)$ by Arzelà-Ascoli’s theorem (Theorem A.5 in Rudin, (1973)), which, with a small abuse of notation, we relabel $\left(u^{n}\right)_{n\in\mathbb{N}}$.
Next, we show (ii). I.e., the above convergent subsequence in $\mathcal{C}^{0}\left(\bar{\mathcal{Y}}\right)$ has a weakly convergent subsequence in ${W}^{1,2,p}_{loc}\left(\mathcal{Y}\right)$. To do so, we prove that $W^{1,2,p}$-norm of the $u^{n}$’s restricted to any compact subset of $\mathcal{Y}$ is eventually uniformly bounded.
@@
$\begin{cases}\max\left\{u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,g-u\right\}=0\text{ in }\mathcal{Y}_{n^{\prime}}\\ u(t,x)=u^{n}(t,x)\text{ on }\partial\mathcal{Y}_{n^{\prime}},\end{cases}$
@@
We then show that there exists a subsequence of $\left(u^{n}\right)_{n\in\mathbb{N}}$ that converges to the limit $u\in\mathcal{C}^{0}\left(\bar{\mathcal{Y}}\right)$ obtained in (i) by a diagonal argument, and, therefore, that $u\in{W}^{1,2,p}_{loc}\left(\mathcal{Y}\right)\cap\mathcal{C}^{0}\left(\bar{\mathcal{Y}}\right)$. For all $n\geq 3$, by (14),
@@
${W}^{1,2,p}\left(\mathcal{Y}_{1}\right)$ is separable and reflexive, and, hence, its closed bounded subsets are weakly sequentially compact by Theorem 1.32 in Demengel et al., (2012). Moreover $W^{1,2,p}(\mathcal{Y}_{1})$ is compactly embedded in $\mathcal{C}^{0}\left(\mathcal{Y}_{1}\right)$ by Rellich-Kondrachov’s theorem (Theorem 2.84 in Demengel et al., (2012)). So, there exists a function $\tilde{u}$ defined on $\bar{\mathcal{Y}}_{T}$ which restriction on $\mathcal{Y}_{1}$ is continuous, and a subsequence $\left(u^{n_{j}}\right)_{j\in\mathbb{N}}\subseteq\left(u^{n}\right)_{n\in\mathbb{N}}$ such that
@@
Proceeding with the diagonal argument, we see that there exists a subsequence $\left(u^{n_{j}}\right)_{j\in\mathbb{N}}\subseteq\left(u_{n}\right)_{n\in\mathbb{N}}$ such that
$\begin{cases}u^{n_{j}}\rightharpoonup u\text{ in }W^{1,2,p}_{loc}\left(\mathcal{Y}\right)\\ u^{n_{j}}\to u\text{ in }\mathcal{C}^{0}\left({\mathcal{Y}}\right),\end{cases}$

### v2

@@ 6.2 Proof of Theorem 1
By Theorem 1.2.2.2 in Grisvard, (2011), $\mathcal{X}$ and $\mathcal{X}_{n}$, $n\in\mathbb{N}$ satisfy a uniform exterior cone condition of size $(L,\theta)$, for some $L,\theta>0$ independent of $n$. In particular $(L,\theta)$ depends only on $M$ and $diam(\mathcal{X})$.
@@
$\begin{cases}\max\left\{u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,g-u\right\}=0\text{ on }\mathcal{Y}_{n}\\ u=b^{n}\text{ on }\partial\mathcal{Y}_{n},\end{cases}$ (Eqn)
@@
By Lemma 3, for all $n\in\mathbb{N}$, there exists a unique $L^{p}$-solution $u^{n}\in W^{1,2,p}\left(\mathcal{Y}_{n}\right)\subset W^{1,2,p}_{loc}\left(\mathcal{Y}_{n}\right)\cap\mathcal{C}^{0}\left(\bar{\mathcal{Y}}_{n}\right)$ of (Eqn). Extending the functions $u^{n}$, $n\in\mathbb{N}$, by letting $u^{n}(t,x)=b(t,x)$ on $\bar{\mathcal{Y}}\setminus\mathcal{Y}_{n}$, we obtain a sequence $\left(u^{n}\right)_{n\in\mathbb{N}}\subset W^{1,2,p}\left(\mathcal{Y}\right)\subset W^{1,2,p}_{loc}\left(\mathcal{Y}\right)\cap\mathcal{C}^{0}\left(\bar{\mathcal{Y}}\right)$.
@@
We start with (i). By Lemma 4 with $D>T\vee diam(\mathcal{X})$ (using Lemma 12.1.9 in Krylov, (2018) to uniformly control $\left\|u^{n}\right\|_{L^{\infty}\left(\mathcal{Y}\right)}$, $n\in\mathbb{N}$), the sequence $\left(u^{n}\right)_{n\in\mathbb{N}}$ is equicontinuous. As a result, it has a convergent subsequence in $\mathcal{C}^{0}\left(\bar{\mathcal{Y}}\right)$ by the Arzelà-Ascoli theorem (Theorem A.5 in Rudin, (1973)), which, with a small abuse of notation, we relabel $\left(u^{n}\right)_{n\in\mathbb{N}}$.
Next, we show (ii), i.e., that the subsequence above, which converges in $\mathcal{C}^{0}\left(\bar{\mathcal{Y}}\right)$, has a weakly convergent subsequence in ${W}^{1,2,p}_{loc}\left(\mathcal{Y}\right)$. To do so, we prove that the $W^{1,2,p}$-norms of the functions $u^{n}$, $n\in\mathbb{N}$, restricted to any given compact subset of $\mathcal{Y}$, are eventually uniformly bounded.
@@
$\begin{cases}\max\left\{u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,g-u\right\}=0\text{ on }\mathcal{Y}_{n^{\prime}}\\ u(t,x)=u^{n}(t,x)\text{ on }\partial\mathcal{Y}_{n^{\prime}},\end{cases}$
@@
We then show that there exists a subsequence of $\left(u^{n}\right)_{n\in\mathbb{N}}$ that converges to the limit $u\in\mathcal{C}^{0}\left(\bar{\mathcal{Y}}\right)$ obtained in (i) by a diagonal argument, and, therefore, that $u\in{W}^{1,2,p}_{loc}\left(\mathcal{Y}\right)\cap\mathcal{C}^{0}\left(\bar{\mathcal{Y}}\right)$. For all $n\geq 3$,(16) implies that
@@
${W}^{1,2,p}\left(\mathcal{Y}_{1}\right)$ is separable and reflexive, and, hence, its closed bounded subsets are weakly sequentially compact by Theorem 1.32 in Demengel et al., (2012). Moreover $W^{1,2,p}(\mathcal{Y}_{1})$ is compactly embedded in $\mathcal{C}^{0}\left(\mathcal{Y}_{1}\right)$ by the Rellich-Kondrachov theorem (Theorem 2.84 in Demengel et al., (2012)). Therefore, there exists a function $\tilde{u}$ defined on $\bar{\mathcal{Y}}_{T}$ which restriction on $\mathcal{Y}_{1}$ is continuous, and a subsequence $\left(u^{n_{j}}\right)_{j\in\mathbb{N}}\subseteq\left(u^{n}\right)_{n\in\mathbb{N}}$ such that
@@
Proceeding with a diagonal argument, we see that there exists a subsequence $\left(u^{n_{j}}\right)_{j\in\mathbb{N}}\subseteq\left(u_{n}\right)_{n\in\mathbb{N}}$ such that
$\begin{cases}u^{n_{j}}\rightharpoonup u\text{ in }W^{1,2,p}_{loc}\left(\mathcal{Y}\right)\\ u^{n_{j}}\to u\text{ in }\mathcal{C}^{0}\left(\bar{\mathcal{Y}}\right),\end{cases}$